OpenStudy (anonymous):

I need some clarification on this problem... please help! :) What is the slope of the line x = -5? Undefined 0 -5 5

jimthompson5910 (jim_thompson5910):

hint: every point on this line has an x coordinate of -5

OpenStudy (anonymous):

So do I have to write an equation?

jimthompson5910 (jim_thompson5910):

what kind of line is this? horizontal? vertical? or diagonal?

OpenStudy (anonymous):

It doesn't say, that's all that's there.. that's what I'm confused about..

jimthompson5910 (jim_thompson5910):

look back at my hint

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

If you plug in any 2 points on this line into a slope formula, what do you get? take a shot...

OpenStudy (anonymous):

Would (-5-1) and (0-(-5) be a good one?

jimthompson5910 (jim_thompson5910):

what are two points that lie on this line?

OpenStudy (anonymous):

I honestly don't know ): Do I just use any coordinate plane? Sorry for all the questions by the way, I'm new to this...

jimthompson5910 (jim_thompson5910):

That's ok, remember every point on this line has an x coordinate of -5

jimthompson5910 (jim_thompson5910):

so two points could be (-5,1) and (-5,2) or (-5,10) & (-5, 15)

jimthompson5910 (jim_thompson5910):

it doesn't matter what the y coordinate is just make sure the x coordinate is -5

jimthompson5910 (jim_thompson5910):

this is all because the equation is x = -5

jimthompson5910 (jim_thompson5910):

once you have your two points, find the slope

OpenStudy (anonymous):

Oh, okay, now I'm starting to understand! I'll figure out what the slope is...

jimthompson5910 (jim_thompson5910):

use: m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

I used yours and got -5 over 10, is that correct?

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

keep in mind subtracting a negative is the same as adding -5-(-5) = -5+5 = ???

OpenStudy (anonymous):

Is that correct? ^^

jimthompson5910 (jim_thompson5910):

no it's not

jimthompson5910 (jim_thompson5910):

you should get 0 in the denominator, try again

jimthompson5910 (jim_thompson5910):

and because you get 0 in the denominator, dividing by zero means the slope is undefined

OpenStudy (anonymous):

Thanks!!!!!!!!!!